# Stochastic Approximation with Block Coordinate Optimal Stepsizes

### v1

@@ Abstract
We consider stochastic approximation with block-coordinate stepsizes and propose adaptive stepsize rules that aim to minimize the expected distance from the next iterate to an optimal point. These stepsize rules employ online estimates of the second moment of the search direction along each block coordinate. The popular Adam algorithm can be interpreted as a particular heuristic for such estimation. By leveraging a simple conditional estimator, we derive a new method that obtains comparable performance as Adam but requires less memory and fewer hyper-parameters. We prove that this family of methods converges almost surely to a small neighborhood of the optimal point, and the radius of the neighborhood depends on the bias and variance of the second-moment estimator. Our analysis relies on a simple aiming condition that assumes neither convexity nor smoothness, thus has broad applicability.
@@ 1 Introduction
where $x\in\mathbf{R}^{n}$ is the decision variable, $\xi$ is a random variable, and $f$ is the loss function. In the context of statistical machine learning (e.g., hastie2009elements, ), $x$ represents the parameters of a prediction model, $\xi$ represents randomly sampled data from some (unknown) probability distribution, and $f(x,\xi)$ is the error in making predictions about $\xi$ using the model parametrized by $x$. We assume that $F$ is bounded below and there exists an optimal solution $x_{*}$.
Suppose that for any pair $x$ and $\xi$, we can evaluate the gradient of $f$ with respect to $x$, denoted as $\nabla f(x,\xi)$. Starting with an initial point $x_{0}\in\mathbf{R}^{n}$, the classical stochastic approximation method RobbinsMonro51 generates a sequence $\{x_{1},x_{2},\ldots\}$ with the update rule
@@
where $\alpha_{t}$ is the stepsize (often called the learning rate in the machine learning literature). The convergence properties of this method are well studied in the stochastic approximation literature (e.g., RobbinsMonro51, ; Wasan69book, ; KushnerYin03book, ). In particular, the stepsizes $\alpha_{t}$ should be nonnegative and unsummable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}=\infty$. Stronger guarantees such as almost sure convergence also require $\alpha_{t}$ to be square-summable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty$ (e.g., RobbinsMonro51, ; Blum1954multidim, ; chung1954, ; ROBBINS1971, ).
Despite the rich literature on their convergence theory, stochastic approximation methods in practice often require heuristics and trial and error in choosing the stepsize sequence $\{\alpha_{t}\}$. Adaptive rules that can adjust stepsizes on the fly have been developed in both the optimization literature (e.g., Kesten58, ; MirzoakhmedovUryasev83, ; RuszczynskiSyski83stats, ; RuszczynskiSyski84IFAC, ; DelyonJuditsky93, ) and by the machine learning community (e.g., Jacobs88DBD, ; Sutton92IDBD, ; Schraudolph1999, ; MahmoodSutton12TuningFree, ). More recently, adaptive algorithms that use coordinate-wise stepsizes have become very popular following the seminal works of AdaGrad DuchiHazanSinger2011adagrad and Adam KingmaBa2014adam , and they demonstrate significantly better performance in training large-scale deep neural networks. In this paper, we propose a family of stochastic approximation methods with block-coordinate stepsizes and present a general framework for their convergence analysis.
@@ 1.1 Stochastic approximation with block-coordinate stepsizes
The Adam algorithm KingmaBa2014adam uses the stochastic momentum in (4) as its search direction and sets the coordinate-wise stepsizes as
@@
Adam KingmaBa2014adam and its variant AdamW Loshchilov2019AdamW (which implements decoupled weight decay) have been very successful in training large-scale deep neural networks. However, theoretical understanding of their superior performance is still incomplete despite many recent efforts (e.g., reddi2019convergence, ; bock2018improvement, ; balles2018dissecting, ; zou2019sufficient, ; defossez2020simple, ; Luo2022neurips, ; kunstner2024heavy, ). On the other hand, there are also many works that propose new variants or alternatives to Adam and AdamW, either starting from fundamental optimization principles (e.g., yao2021adahessian, ; gomes2024adafisher, ; hwang2024fadam, ; lin2024can, ; jordan2024muon, ; pethick2025LMOs, ) or based on empirical algorithmic search (e.g., chen2023symbolic, ; zhang2024adam, ). But all have limited success. Adam and especially AdamW are still the most widely used algorithms for training large deep learning models, and their effectiveness remains a myth.
@@ 1.2 Contributions and outline
  • In Section 2, we derive BCOS by minimizing the expected distance from the next iterate to an optimal point. While the optimal stepsizes cannot be computed exactly, we make several simplifications and approximate the second moments of the search direction with simple EMA estimators.
  • In Section 3, we instantiate BCOS with specific search directions. In particular, we show that RMSProp rmsprop2012lecture and Adam KingmaBa2014adam can be interpreted as special cases of BCOS. With the stochastic momentum as search direction, we leverage a simple conditional estimator to derive a variant that requires fewer optimizer states and hyper-parameters than Adam. Integrating with decoupled weight decay Loshchilov2019AdamW gives the BCOSW variants that enjoy comparable performance as AdamW.
@@
The general form of stochastic approximation (e.g., RobbinsMonro51, ; Wasan69book, ; KushnerYin03book, ) concerns solving a nonlinear system of equations $G(x)=0$ where the map $G:\mathbf{R}^{n}\to\mathbf{R}^{n}$ only admits noisy evaluations in the form of $g(\cdot,\xi)$, where $\xi$ is a random variable, under the assumption that $G(x)=\mathbf{E}_{\xi}[g(x,\xi)]$. In this case, stochastic approximation takes the form
@@ 1.3 Notations
[…]
The $\textup{sign}(\cdot)$ function is defined as $\textup{sign}(\alpha)=1$ if $\alpha>0$, $-1$ if $\alpha<0$ and 00 if $\alpha=0$.
@@ 2.1 Block-coordinate optimal stepsizes
We consider the distance of the next iterate $x_{t+1}$ to an optimal point $x_{*}$. Specifically, in terms of the squared Euclidean distance,
@@ 2.2 Simplification and approximation
We first try to avoid the direct reliance on $x_{*}$. To this end, rewrite the numerator in (10) as
@@
Notice that any $\alpha_{t,k}$ we choose in practice may only be a (very rough) approximation of $\|x_{t,k}-x_{*,k}\|\cos\theta_{t,k}$. In particular, while the optimal stepsizes $\widehat{\gamma}_{t,k}$ can be positive or negative, it is very hard to even correctly estimate the sign of the inner product $\langle x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\rangle$. We take the pragmatic approach of restricting $\alpha_{t,k}>0$, effectively being optimistic that the expected direction $-\mathbf{E}_{t}[d_{t,k}]$ is always pointing towards $x_{*,k}$ (i.e., assuming $\cos\theta_{t,k}>0$).
A further simplification is to use a common stepsize schedule $\alpha_{t}$ across all blocks, i.e., $\alpha_{t,k}=\alpha_{t}$ for all $k=1,\ldots,m$. This can be a reasonable choice for training deep neural networks, where the model parameters are initialized randomly coordinate-wise such that $\mathbf{E}[\|x_{0,k}\|]$ is constant for each coordinate $k$ HeZhangRenSun2015 ; RobertsYaidaHanin2022 ; DinanYaidaZhang2023 . This brings us to
@@
We note that with some abuse of notation, here $\alpha_{t}$ denotes a scalar, not a vector of $(\alpha_{t,1},\ldots,\alpha_{t,k})$. This simplification motivates us to link $\alpha_{t}$ to the average of $\alpha_{t,k}$ in (11) and further relate to the overall distance $\|x_{t}-x_{*}\|$. Therefore, we expect $\alpha_{t}$ to decrease as $\|x_{t}-x_{*}\|$ gradually shrinks. A simple strategy is to use a monotonic stepsize schedule on $\alpha_{t}$, such as the popular cosine decay (loshchilov2016sgdr, ) or linear decay (defazio2024optimallineardecaylearning, ).
@@ 2.3 BCOS with one EMA estimator
Apparently we have $\rho_{t,k}\in[0,1]$, and the two boundary values of 00 and $1$ correspond to the cases of zero signal and full signal (zero noise) respectively.
@@
are invariant of scaling of the search directions $d_{t,k}$, which makes the tuning of $\alpha^{\prime}_{t}$ much easier in practice. This invariance property is also shared by Adam KingmaBa2014adam , as seen from (5) and (6). The similarity of Adam and the simplified BCOS in (17) is apparent, and we will explain their connection in detail in the next section.
@@ 3.1 BCOS with EMA estimator
We immediately recognize that BCOS-g is exactly the RMSprop algorithm rmsprop2012lecture , which is one of the first effective algorithms for training deep neural networks. Our BCOS framework gives a novel interpretation of RMSprop and its effectiveness. In the special case with $\beta=0$ and $\epsilon=0$, we have $v_{t}=g_{t}^{2}$, and BCOS-g becomes the sign gradient method
@@
which also received significant attention in the literature Polyak1973pseudo ; bernstein2018signsgd ; safaryan2021signdescent ; jiang2024efficient .
@@
Using the stochastic momentum as search direction has a long history in stochastic approximation (e.g., GupalBazhenov72, ; Polyak77comparison, ; RuszczynskiSyski83stats, ). It has become the default option for modern deep learning due to its superior performance compared with using plain stochastic gradients. In Algorithm 2, we use $m_{t}$ to denote the momentum, which is a standard notation in the machine learning literature. We call it BCOS-m to signify the use of momentum as the search direction. BCOS-m employs a second smoothing factor $\beta_{2}$ to calculate the EMA of $m_{t}^{2}$. The two smoothing factors $\beta_{1}$ and $\beta_{2}$ do not need to be the same and can be chosen independently in practice. Similar to BCOS-g, the special case of BCOS-m with $\beta_{2}=0$ and $\epsilon=0$ corresponds to the sign-momentum method
@@
In both BCOS-g and BCOS-m, we initialize $v_{-1}=g_{0}^{2}$, therefore $v_{0}=g_{0}^{2}$ as well. This avoids the initial bias caused by setting $v_{-1}=0$. An alternative way for bias correction is to initialize $m_{-1}$ and $v_{-1}$ as the zero vector, and then multiply $m_{t}$ and $v_{t}$ by $1/(1-\beta^{t+1})$, which quickly converges to $1$, as done in Adam KingmaBa2014adam . Our numerical experiments show that these two choices do not have significant differences in practice.
@@ 3.2 BCOS with conditional estimator
Recall that the optimal stepsizes $\widehat{\gamma}_{t,k}$ in (10) and their simplifications $\widetilde{\gamma}_{t,k}$ in (13) are all based on conditional expectation. In Section 3.1, we used coordinate-wise EMA of $d_{t}^{2}$ to approximate the conditional expectation $\mathbf{E}_{t}[d_{t}^{2}]$, i.e., $v_{t}$ as estimator of $\mathbf{E}_{t}[d_{t}^{2}]$ in BCOS-g and of $\mathbf{E}_{t}[m_{t}^{2}]$ in BCOS-m, respectively. In this section, we show that with $m_{t}$ as the search direction, we can exploit its update form to derive effective conditional estimators that can avoid using EMA.
@@
input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\epsilon>0$
@@
     $v_{t}=\bigl{(}1-(1-\beta)^{2}\bigr{)}m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}$
@@
Algorithm 4 BCOSW-c
input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\lambda\geq 0$, $\epsilon>0$
$m_{-1}=\nabla f(x_{0},\xi_{0})$
for $t=0,1,2,\ldots$ do
     $g_{t}=\nabla f(x_{t},\xi_{t})$
     $m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$
     $v_{t}=\bigl{(}1-(1-\beta)^{2}\bigr{)}m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}$
     $x_{t+1}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha_{t}\lambda)}x_{t}-\alpha_{t}\frac{m_{t}}{\sqrt{v_{t}}+\epsilon}$
$m_{t}=\beta m_{t-1}+(1-\beta)g_{t}.$ (23)
To derive an estimator of $\mathbf{E}_{t}[m_{t}^{2}]$, we expand the square and take the expectation of each term:
$=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}\odot\mathbf{E}_{t}[g_{t}]+(1-\beta)^{2}\mathbf{E}_{t}\bigl{[}g_{t}^{2}\bigr{]}.$ (24)
Our numerical experiments confirmed that this indeed is a good estimator which obtains performance comparable to Adam.
While effective in practice, the estimator in (25) looks complicated and does not carry much intuition. We are motivated to derive another one that is much simpler and as effective. The key is to approximate $\mathbf{E}[g_{t}]$ in (24) with $m_{t-1}$ instead of $m_{t}$, which results in
@@
This estimator is much simpler than (25). Moreover, it resembles the standard EMA estimator in Adam, shown in (6), with an effective smoothing factor
$\beta^{\prime}=1-(1-\beta)^{2},$
but with $v_{t-1}$ replaced by $m^{2}_{t-1}$. As a result, the estimator in (26) does not need to store $v_{t-1}$, thus requiring fewer optimizer states (less memory). This also explains that the second smoothing factor in Adam, $\beta_{2}$, corresponding to $\beta^{\prime}$ here, should be much larger or closer to 1 than $\beta$. Specifically, $\beta=0.9$ roughly corresponds to $\beta^{\prime}=0.99$. The estimator in (26) eliminates $\beta_{2}$ as a second hyper-parameter.
Finally, replacing $v_{t}$ in BCOS-m with the one in (26) produces Algorithm 3. We call it BCOS-c to signify the conditional estimator. It has fewer optimizer states and fewer hyper-parameters to tune than BCOS-m and Adam. We note the slight abuse of notation in using the symbol $v_{t}$ for both the EMA estimator in BCOS-g and -m and the conditional estimator here. In general, $v_{t}$ denotes an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$. The corresponding choice of $d_{t}$ and type of estimator should be clear from the context.

### v2

@@ Abstract
We consider stochastic approximation with block-coordinate stepsizes and propose adaptive stepsize rules that aim to minimize the expected distance from the next iterate to an (unknown) target point. These stepsize rules employ online estimates of the second moment of the search direction along each block coordinate. The popular Adam algorithm can be interpreted as a variant with a specific estimator. By leveraging a simple conditional estimator, we derive a new method that obtains competitive performance against Adam but requires less memory and fewer hyper-parameters. We prove that this family of methods converges almost surely to a small neighborhood of the target point, and the radius of the neighborhood depends on the bias and variance of the second-moment estimator. Our analysis relies on a simple aiming condition that assumes neither convexity nor smoothness, thus has broad applicability.
Keywords: stochastic approximation, block coordinate methods, conditional estimator, aiming condition, almost sure convergence.
@@ 1 Introduction
where $x\in\mathbf{R}^{n}$ is the decision variable, $\xi$ is a random variable, and $f$ is the loss function. In the context of statistical machine learning (e.g., hastie2009elements), $x$ represents the parameters of a prediction model, $\xi$ represents randomly sampled data from some (unknown) probability distribution, and $f(x,\xi)$ is the error in making predictions about $\xi$ using the model parametrized by $x$. We assume that $F$ is bounded below.
Suppose that for any pair $x$ and $\xi$, we can evaluate the gradient of $f$ with respect to $x$, denoted as $\nabla f(x,\xi)$. Starting with an initial point $x_{0}\in\mathbf{R}^{n}$, the classical stochastic approximation method (RobbinsMonro51) generates a sequence $\{x_{1},x_{2},\ldots\}$ with the update rule
@@
where $\alpha_{t}$ is the stepsize (often called the learning rate in the machine learning literature). The convergence properties of this method are well studied in the stochastic approximation literature (e.g., RobbinsMonro51; Wasan69book; KushnerYin03book). In particular, the stepsizes $\alpha_{t}$ should be nonnegative and unsummable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}=\infty$. Stronger guarantees such as almost sure convergence also require $\alpha_{t}$ to be square-summable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty$ (e.g., RobbinsMonro51; Blum1954multidim; chung1954; ROBBINS1971).
Despite the rich literature on their convergence theory, stochastic approximation methods in practice often require heuristics and trial and error in choosing the stepsize sequence $\{\alpha_{t}\}$. Adaptive rules that can adjust stepsizes on the fly have been developed in both the optimization literature (e.g., Kesten58; MirzoakhmedovUryasev83; RuszczynskiSyski83stats; RuszczynskiSyski84IFAC; DelyonJuditsky93) and by the machine learning community (e.g., Jacobs88DBD; Sutton92IDBD; Schraudolph1999; MahmoodSutton12TuningFree). More recently, adaptive algorithms that use coordinate-wise stepsizes have become very popular following the seminal works of AdaGrad (DuchiHazanSinger2011adagrad) and Adam (KingmaBa2014adam), and they demonstrate significantly better performance in training large-scale deep neural networks.
In this paper, we propose a family of stochastic approximation methods with block-coordinate stepsizes and present a general framework for their convergence analysis.
@@ 1.1 Stochastic approximation with block-coordinate stepsizes
The Adam algorithm (KingmaBa2014adam) uses the stochastic momentum in (4) as its search direction and sets the coordinate-wise stepsizes as
@@
Adam (KingmaBa2014adam) and its variant AdamW (Loshchilov2019AdamW), which implements decoupled weight decay, have been very successful in training large-scale deep neural networks. However, theoretical understanding of their superior performance is still incomplete despite many recent efforts (e.g., bock2018improvement; balles2018dissecting; reddi2019convergence; zou2019sufficient; defossez2020simple; Luo2022neurips; kunstner2024heavy). On the other hand, there are also many works that propose new variants or alternatives to Adam and AdamW, either starting from fundamental optimization principles (e.g., yao2021adahessian; gomes2024adafisher; hwang2024fadam; jordan2024muon; lin2024can; pethick2025LMOs) or based on empirical algorithmic search (e.g., chen2023symbolic; zhang2024adam). But all have limited success. Adam and especially AdamW are still the most widely used algorithms for training large deep learning models, and their effectiveness remains a myth.
@@ 1.2 Contributions and outline
  • In Section 2, we derive BCOS by minimizing the expected distance from the next iterate to a target point. While the optimal stepsizes cannot be computed exactly, we make several simplifications and approximate the second moments of the search direction with simple EMA estimators.
  • In Section 3, we instantiate BCOS with specific search directions. In particular, we show that RMSProp (rmsprop2012lecture) and Adam (KingmaBa2014adam) can be interpreted as special cases of BCOS. With the stochastic momentum as search direction, we leverage a simple conditional estimator to derive a variant that requires less memory and fewer hyper-parameters than Adam. Integrating with decoupled weight decay (Loshchilov2019AdamW) gives the BCOSW variants that enjoy competitive performance as AdamW.
@@
The general form of stochastic approximation (e.g., RobbinsMonro51; Wasan69book; KushnerYin03book) concerns solving a nonlinear system of equations $G(x)=0$ where the map $G:\mathbf{R}^{n}\to\mathbf{R}^{n}$ only admits noisy evaluations in the form of $g(\cdot,\xi)$, where $\xi$ is a random variable, under the assumption that $G(x)=\mathbf{E}_{\xi}[g(x,\xi)]$. In this case, stochastic approximation takes the form
@@ 1.3 Notations
Throughout this paper, $\langle\cdot,\cdot\rangle$ denotes the standard inner product in the Euclidean space $\mathbf{R}^{n}$ and $\|\cdot\|$ denotes the induced Euclidean norm. The $\textup{sign}(\cdot)$ function is defined as $\textup{sign}(\alpha)=1$ if $\alpha>0$, 1-1 if $\alpha<0$ and 0 if $\alpha=0$; for vectors, it applies element-wise.
@@ 2.1 Block-coordinate optimal stepsizes
We consider the distance of the next iterate $x_{t+1}$ to an (unknown) target point $x_{*}$. Specifically, in terms of the squared Euclidean distance,
@@ 2.2 Simplification and approximation
We first try to avoid the direct reliance on $x_{*}$, which is unknown to us. To this end, rewrite the numerator in (10) as
@@
Notice that any $\alpha_{t,k}$ we choose in practice may only be a (very rough) approximation of $\|x_{t,k}-x_{*,k}\|\cos\theta_{t,k}$. In particular, while the optimal stepsizes $\widehat{\gamma}_{t,k}$ can be positive or negative, it is very hard to even correctly estimate the sign of the inner product $\langle x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\rangle$. We take the pragmatic approach of restricting $\alpha_{t,k}>0$, effectively being optimistic that the expected direction $\mathbf{E}_{t}[d_{t,k}]$ is always pointing towards $x_{*,k}$ (i.e., assuming $\cos\theta_{t,k}>0$).
A further simplification is to use a common stepsize schedule $\alpha_{t}$ across all blocks, i.e., $\alpha_{t,k}=\alpha_{t}$ for all $k=1,\ldots,m$. This can be a reasonable choice for training deep neural networks, where the model parameters are initialized randomly coordinate-wise such that $\mathbf{E}[\|x_{0,k}\|]$ is constant for each coordinate $k$ (HeZhangRenSun2015; RobertsYaidaHanin2022; DinanYaidaZhang2023). This brings us to
@@
We note that with some abuse of notation, here $\alpha_{t}$ denotes a scalar, not a vector of $(\alpha_{t,1},\ldots,\alpha_{t,k})$. This simplification motivates us to link $\alpha_{t}$ to the average of $\alpha_{t,k}$ in (11) and further relate to the overall distance $\|x_{t}-x_{*}\|$. Therefore, we expect $\alpha_{t}$ to decrease as $\|x_{t}-x_{*}\|$ gradually shrinks. A simple strategy is to use a monotonic stepsize schedule on $\alpha_{t}$, such as the popular cosine decay (loshchilov2016sgdr) or linear decay (defazio2024optimallineardecaylearning).
@@ 2.3 BCOS with one EMA estimator
Apparently we have $\rho_{t,k}\in[0,1]$, and the two boundary values of 0 and $1$ correspond to the cases of zero signal and full signal (zero noise) respectively.
@@
are invariant of scaling of the search directions $d_{t,k}$, which makes the tuning of $\alpha^{\prime}_{t}$ much easier in practice. This invariance property is also shared by Adam (KingmaBa2014adam), as seen from (5) and (6). The similarity of Adam and the simplified BCOS in (17) is apparent, and we will explain their connection in detail in the next section.
@@ 3.1 BCOS with EMA estimator
We immediately recognize that BCOS-g is exactly the RMSprop algorithm (rmsprop2012lecture), which is one of the first effective algorithms for training deep neural networks. Our BCOS framework gives a novel interpretation of RMSprop and its effectiveness. In the special case with $\beta=0$ and $\epsilon=0$, we have $v_{t}=g_{t}^{2}$, and BCOS-g becomes the sign gradient method
@@
which also received significant attention in the literature (Polyak1973pseudo; bernstein2018signsgd; safaryan2021signdescent; jiang2024efficient).
@@
Using the stochastic momentum as search direction has a long history in stochastic approximation (e.g., GupalBazhenov72; Polyak77comparison; RuszczynskiSyski83stats). It has become the default option for modern deep learning due to its superior performance compared with using plain stochastic gradients. In Algorithm 2, we use $m_{t}$ to denote the momentum, which is a standard notation in the machine learning literature. We call it BCOS-m to signify the use of momentum as the search direction. BCOS-m employs a second smoothing factor $\beta_{2}$ to calculate the EMA of $m_{t}^{2}$. The two smoothing factors $\beta_{1}$ and $\beta_{2}$ do not need to be the same and can be chosen independently in practice. Similar to BCOS-g, the special case of BCOS-m with $\beta_{2}=0$ and $\epsilon=0$ corresponds to the sign-momentum method
@@
In both BCOS-g and BCOS-m, we initialize $v_{-1}=g_{0}^{2}$, therefore $v_{0}=g_{0}^{2}$ as well. This avoids the initial bias caused by setting $v_{-1}=0$. An alternative way for bias correction is to initialize $m_{-1}$ and $v_{-1}$ as the zero vector, and then multiply $m_{t}$ and $v_{t}$ by $1/(1-\beta^{t+1})$, which quickly converges to $1$, as done in Adam (KingmaBa2014adam). Our numerical experiments show that these two choices do not have significant differences in practice.
@@ 3.2 BCOS with conditional estimator
Recall that the derivations of optimal stepsizes $\widehat{\gamma}_{t,k}$ and $\widetilde{\gamma}_{t,k}$ in Section 2 are all based on conditional expectation $\mathbf{E}_{t}[\cdot]$. In Section 3.1, we used coordinate-wise EMA to approximate the conditional expectations, specifically, $v_{t}$ as an estimator of $\mathbf{E}_{t}[d_{t}^{2}]$ in BCOS-g and an estimator of $\mathbf{E}_{t}[m_{t}^{2}]$ in BCOS-m. In this section, we show that with $m_{t}$ as the search direction, we can exploit its update form to derive more effective conditional estimators.
$m_{t}=\beta m_{t-1}+(1-\beta)g_{t}.$ (23)
To derive an estimator of $\mathbf{E}_{t}[m_{t}^{2}]$, we expand the square and take the conditional expectation of each term:
$=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}\odot\mathbf{E}_{t}[g_{t}]+(1-\beta)^{2}\mathbf{E}_{t}\bigl[g_{t}^{2}\bigr],$ (24)
The estimator in (25) does not need to store $v_{t-1}$, thus requiring less memory (fewer optimizer states). Moreover, this estimator eliminates $\beta_{2}$ as a second hyper-parameter.
@@
input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\epsilon>0$,
@@
  $v_{t}=\beta^{2}m^{2}_{t-1}+2\beta(1-\beta)m_{t-1}\odot m_{t}+(1-\beta)^{2}g^{2}_{t}$
  ($\Big$(Simple alternative: $v_{t}=\bigl(1-(1-\beta)^{2}\bigr)m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}$ $\Big)$
@@
Replacing $v_{t}$ in BCOS-m with (25) leads to Algorithm 3. We call it BCOS-c to signify the conditional estimator.
The estimator in (25) can be further simplified if we approximate $\mathbf{E}_{t}[g_{t}]$ in (24) with $m_{t-1}$ instead of $m_{t}$, which results in
@@
This estimator resembles the standard EMA estimator in Adam, shown in (6), with an effective smoothing factor
$\beta^{\prime}=1-(1-\beta)^{2},$ (27)
but with $v_{t-1}$ replaced by $m^{2}_{t-1}$. This also explains that the second smoothing factor in Adam, $\beta_{2}$, corresponding to $\beta^{\prime}$ here, should be much larger or closer to 1 than $\beta$. Specifically, $\beta=0.9$ roughly corresponds to $\beta^{\prime}=0.99$.
We include the estimator (26) as a simple alternative of (25) in Algorithm 3. Empirically we found that it behaves almost identically as (25) on the experiments we conduct in Section 4. Concurrently, zhang2025adams propose a similar algorithm with
$v_{t}=\beta^{\prime}m_{t-1}^{2}+(1-\beta^{\prime})g_{t}^{2},$
and show that tuning $\beta^{\prime}$ as a hyper-parameter can obtain slightly better performance than using the value in (27).
Algorithm 4 BCOSW{-g, -m, -c}
input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\epsilon>0$, $\lambda\geq 0$
for $t=0,1,2,\ldots$ do
  compute $d_{t}$ ($g_{t}$ or $m_{t}$) and $v_{t}$ as in Algorithm 1, 2 or 3 (-g, -m or -c, respectively)
  $x_{t+1}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha_{t}\lambda)}x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}$